OpenStudy (anonymous):

7. Solve the equation 5/4x-4 - 6/x+1 = -1 Help Please!!!

OpenStudy (johnweldon1993):

\[\large \frac{5}{4x - 4} - \frac{6}{x+1} = -1\] like that?

OpenStudy (anonymous):

yes!!!!! please help

OpenStudy (anonymous):

hello?